OpenStudy (anonymous):

How do I set this up? If 3 roses and 2 tulips cost $1.50 more than 2 roses and 3 tulips then how much more does a rose cost than a tulip? Thanks

OpenStudy (bibby):

decide on variables. we can use r and t if 3r + 2t = 2r + 3t + 1.50 can you rewrite this by solving for r or t?

OpenStudy (bibby):

this actually might not be right. I'm trying to get it all in terms of one variable. keep bumping your question

OpenStudy (anonymous):

ok The teacher gives us problems that we have not studies. We have not solved for variables yet but I know how to do it sort of, It is an Ice Storm day question.

OpenStudy (bibby):

This might be unreadable but I think I got the price of a rose. Let me know if you want me to redo anything or go into more detail

OpenStudy (ranga):

3r + 2t = 2r + 3t + 1.50 3r - 2r + 2t - 3t = 1.50 r - t = 1.50 The answer is $1.50 because they are asking for r - t